OpenStudy (anonymous):

June has a set of folding chairs. If she arranges the chairs in 5 rows, she has 2 chairs left over. If she arranges them in 3 rows of the same length, she has 14 left over. How many chairs does she have?

OpenStudy (anonymous):

I think that would just be setting up two equations like so: let S = set of chairs, R = row length. \[5R + 2 = S\]\[3R +14 = S\]and then solving for S.

OpenStudy (anonymous):

i forgot to do the second step...HAHA! s = 6, but it asked how many chairs there are. 6x5+2 = 32 chairs...duh. thank you.